# Defects, parcellation, and renormalized negative diffusivities in non-homogeneous oscillatory media

### v1

@@ Abstract
Coupling among oscillators in spatially-extended systems tends to lock their frequency at a common value. In the presence of spatial non-homogeneities, locking of different regions at different frequencies leads to parcellation, i.e., a series of synchronized clusters (plateaus). Motivated by rhythmic dynamics in physiological systems, we consider a Ginzburg-Landau (GL) model with a gradient of natural frequencies. We determine the scaling of the number of plateaus and their typical length vs dynamical parameters. Plateaus are separated by defects, where the amplitude of the GL field vanishes and phase differences are reset. For Dirichlet boundary conditions, we use asymptotic methods to determine the field profile around defects. For Neumann boundary conditions, we relate the stability phase diagram and defects’ precursors to the spectrum of the non-Hermitian Bloch-Torrey equation, originally introduced for nuclear magnetic resonance. In the non-linear regime, we trace the formation of defects to a non-linear renormalization of the diffusivity, which leads to spatially-modulated negative values and an instability that drives amplitude modulation.
@@
Spatially-extended oscillations appear in a variety of situations, ranging from engineering to living, physical, and social systems Kuramoto (1984); Pikovsky (2001); Strogatz (2004). Hallmarks of nonlinearities and spatial couplings that characterize those systems are synchronization among different oscillators, and entrainment by external forcing. Synchronization is the oscillation in unison of oscillators that in isolation have different frequencies yet lock because of their mutual coupling. A spectacular example is the synchronized flashing of thousands of firefly males during mating season Buck (1966); Peleg (2020). Other well-studied examples include locomotion, especially by fish Grillner (2020), and electrical waves in neural computations Ermentrout (2001); Muller (2018). Entrainment by external forcing is the locking of oscillators to the driving rhythm. A physiological instance that impacts functional magnetic resonant imaging (fMRI) is vasomotion Haddock (2005); Drew (2020). Arterioles that source blood to the brain have the diameter of their walls driven by oscillations of the surrounding smooth muscle cells. In the presence of external neuronal stimuli, a strong vasomotion increase is observed at the drive frequency Broggini (2024).
@@
A number of theoretical methods have been developed to grasp the above phenomena. Phase models Kuramoto (1984); Pikovsky (2001) are rooted in the insight that weak couplings among oscillators perturb the stable limit cycle of each oscillator in a specific manner: Transverse perturbations relax rapidly, while phase perturbations along the cycle can be large and do not relax (zero modes). Technically, the limit cycle and its isochrones Winfree (2001) are unperturbed, while weak perturbations drive the phase. The major simplification is that each oscillator is described by only a single variable, its phase in the limit cycle, rather than the original, generally multidimensional, description (see Supplementary Material NoteSM for their relationship). The famous Kuramoto transition to synchronization Kuramoto (1984) was obtained within this framework.
@@
We focus on spatially non-homogeneous oscillatory media. Non-homogeneity may be caused by spatial variations of the parameters, e.g., the frequency of oscillators, or non-uniform external drive. Our motivation stems from two natural systems under intense experimental investigation : the aforementioned cortical vasomotion Haddock (2005); Drew (2020); Broggini (2024), and peristalsis Diamant (1969); Parsons (2020). Interstitial cells in the small intestine of mammals bear spontaneous oscillations, which are synchronized by gap junction couplings, and generate rhythmic waves of contraction. Non-homogeneity is brought by the frequencies of the oscillators in isolation decreasing aborally in a roughly linear fashion from the stomach to the colon (see Fig. 1a), found by post hoc sectioning of the intestine Diamant (1969). In the intact intestine, couplings result in parcellation, i.e., a series of frequency plateaus as in Fig. 1a. Oscillations within the plateaus are regular, while their boundary displays “waxing and waning” (see Fig. 1b) that results from the interference of mismatched adjacent cell populations Diamant (1969).
@@
Figure 1: (Color online) Comparison between experimental peristalsis and our GL modeling data. (a) Average frequencies recorded in isolated segments of cat intestine (crosses ; the black dashed line is their linear fit) and in the intact intestine (circles); (b) Time-course signals in the bulk of two distinct plateaus (top and bottom), where relatively regular oscillations are observed, and in the region of “waxing and waning” in between. Data were extracted from figures in Diamant (1969) as described in NoteSM . Panels (a) and (b) refer to two different experimental realizations so we do not associate the curves in (b) to specific points in (a). In panels (c) and (d) we show the same curves as in (a) and (b) obtained from our GL modeling with the use of the procedure detailed in NoteSM . The plateau architecture is captured, as well as the structure of the three time courses that we locate by their colors. The inset shows the full simulation domain.
@@
To understand parcellation, it is natural to consider a chain of weakly coupled oscillators with a gradient of natural frequencies. Ref. Ermentrout (1984) pioneered this approach and demonstrated the existence of plateaus. However, as already noted in Ermentrout (1986), the strong amplitude modulation at the boundary of plateaus is incompatible with the weak coupling approximation. A first alternative is to consider chains of nonlinear oscillators, viz., van der Pol type Osipov (1998); Menzel (2011), which is relevant for density waves in dusty plasmas Menzel (2011). Numerical simulations confirmed the existence of plateaus and evidenced “oscillator death” for large enough gradients, i.e., the presence of regions with oscillators unexcited Ermentrout (1990); Reddy (2000); Ozden (2004). A second alternative is to consider a Ginzburg-Landau (GL) model (see Aranson (2002)) with a gradient of natural frequencies. The advantage is that the GL equation is the universal normal form close to a supercritical Hopf bifurcation. Oscillations are described by a complex field with a phase as well as an amplitude, contrary to phase models. Refs. Ermentrout (1986); Akopov (2003); Anishchenko (2005) numerically demonstrated plateaus in the GL equation and analyzed its full locking and oscillator death at large gradients. In the GL formulation, the “waxing and waning” at the boundary of plateaus is associated to defects, i.e., locations where the amplitude of the complex field vanishes at specific times and its phase is thus undefined Pikovsky (2001).
Here, we shall revisit the GL equation with a gradient of natural frequencies to provide insight into the formation and properties of plateaus, as well as the dynamics at high gradients. Specifically, we determine the scaling of length and number of plateaus, and calculate by asymptotic analysis the field profile in the region around a defect. For strong gradients, we identify the stability chart of the oscillator death state. In the nonlinear regime, we derive that non-linear renormalization of the diffusivity yields negative values that destabilize full locking and drive amplitude modulation and defects. Finally, we show that the experimental curves in Fig. 1 are captured by a GL model with an additional advective term.
@@
We consider a one-dimensional Ginzburg-Landau (GL) equation for the complex field $A(x,t)$ :
@@
which is the Stuart-Landau normal form for oscillators close to a supercritical Hopf bifucation, with diffusion accounting for spatial local couplings Kuramoto (1984); Pikovsky (2001). We rescaled the amplitude of $A$ to match the cubic and the instability coefficient, $r>0$, which we take real, along with taking $D$ as real and positive. These assignments are justified a posteriori by the agreement with data in Fig. 1. Note that the cubic nonlinearity is dictated by gauge invariance under $A\mapsto Ae^{i\phi}$. The linear gradient $\omega(x)=gx$ introduces spatial non-homogeneity. It also breaks parity, which allows a priori for an additional term $\propto\nabla A$ that we shall consider later.
A straightforward means to demonstrate parcellation is to take Dirichlet boundary conditions $A\left(x=\pm L/2,t\right)=e^{\pm i\frac{gL}{2}t}$ at the boundaries of the interval $[-L/2,L/2]$. The symmetries $\mathrm{Re}[A(x)]=\mathrm{Re}[A(-x)]$ and $\mathrm{Im}[A(x)]=-\mathrm{Im}[A(-x)]$ allow us to double the resolution (see End Matter) and impose $\mathrm{Im}[A(0)]=0$. The average frequency at each spatial point shown in Figs. 2b,c confirms the presence of extended regions of synchronization. Note that the equality of average frequencies ensures that phases of oscillators in a plateau vary in time with the same dominant term but may not keep a fixed phase difference (strict phase locking) Pikovsky (2001).
The simplest example of parcellation has two plateaus symmetric around the origin, locked at the boundary frequencies $\mp Lg/2$, as shown in Fig. 2c. The jump between the two plateaus dictates the characteristic time $2\pi/gL$ for accumulation of phase differences and amplitude modulation. When the phases in the two plateaus reach $-\pi$ and $+\pi$, a defect appears at $x=0$. The amplitude oscillates and reaches $A=0$ at the formation of the defect, when the phase is undefined Pikovsky (2001), the $2\pi$ difference is reset and the process restarts (see Movie 1 in Ref. NoteSM ). To address amplitude modulation, we decompose $A(x,t)=a(x,t)e^{i\varphi(x,t)}$ and rewrite Eq. (1) as :
@@
to identify $-Da(\nabla\varphi)^{2}$ as responsible for damping $a$ as $|\nabla\varphi|$ increases, eventually driving $a=0$ at the defect.
An estimate of the spatial extension of amplitude modulation is calculated using a WKB expansion of a field profile with two plateaus, as shown in Fig. 2c. WKB is the classical method for small diffusion limits Bender (1999). Three non-dimensional numbers will be useful to understand the dynamics of Eq. (1) :
$\alpha\equiv\frac{L^{3}|g|}{D}\,;\quad\beta\equiv\frac{Dg^{2}}{r^{3}}\,;\quad\gamma\equiv\left(\beta\alpha^{4}\right)^{\frac{1}{3}}=\frac{L^{4}g^{2}}{rD}\,.$ (4)
The first yields an estimate of the ratio between the gradient and the diffusive terms, with $\alpha\gg 1$ corresponding to the small diffusion limit. Movie 1 in NoteSM shows that the field can be reduced to a quasi-static profile with an oscillatory region around $x=0$ (location of the defect) that shrinks as $|g|$ grows. We found that the WKB profile vs $\tilde{x}\equiv x+L/2$ in the bulk of the left plateau for large $\alpha$ and $\beta$ reads (see Figs. 2d,e and End Matter) :
$A\propto\frac{e^{-\frac{itgL}{2}-\sqrt{\frac{2}{9}}\left(1-i\,\rm{sgn}\left(g\right)\right)\left(\frac{\tilde{x}}{\ell_{p}}\right)^{3/2}}}{\tilde{x}^{1/4}}\,;\;\ell_{p}\equiv\left(\frac{D}{|g|}\right)^{1/3}\,.$ (5)
Eq. (5) is generalized to moderate $\beta$’s in NoteSM .
Figure 2: (Color online) Descriptive properties of plateaus. Panel (a) shows the number of plateaus $N_{p}$ vs $\alpha^{1/3}$ for Dirichlet (circles) and Neumann (squares) boundary conditions for $\beta=10^{-2}$ (empty symbols) and $10^{-4}$ (full symbols) ; the insets show plateaus for Dirichlet boundary conditions with $L=1$, $\beta=10^{-4}$ and (b) $\alpha^{1/3}=21.56$ or (c) $\alpha^{1/3}=2.71$. Panels (d) and (e) illustrate amplitude and phase for $\alpha=4\times 10^{4}$ and $\beta=1$ (red) ; $\alpha=2\times 10^{5}$ and $\beta=10$ (green) ; $\alpha=1.6\times 10^{5}$ and $\beta=1000$ (blue). The agreement of our WKB prediction (black dashed lines) in the half-interval $[-L/2,0]$ confirms its validity for large $\alpha$ and moderate/large $\beta$.
Figure 3: Eigenvalue spectra and phase space portrait. Panels (a)-(b) report the opposite of the real and the imaginary part of the first six eigenvalues $\sigma_{n}$ vs $\alpha^{1/3}$ for the Bloch-Torrey Eq. (7) (ordered by their real part). The imaginary part vanishes before the branching point while two complex conjugate eigenvalues per branch are present afterwards. (c) Phase space portrait. Stable: all eigenvalues have negative real part. Locking: all eigenvalues are real and at least one has positive real part. A locked state is observed in the nonlinear regime Ermentrout (1986). Nonlinear Locking: locking is observed in the nonlinear regime in spite of one or more complex pairs of eigenvalues (linear precursors of plateaus) being present. Plateaus: at least a complex pair of eigenvalues is present and the nonlinear regime features multiple plateaus; the dashed line is the asymptotics $\beta=\gamma_{c}^{3}\alpha^{-4}$ for the critical curve separating locking from multiple plateaus, with the constant $\gamma_{c}$ predicted by Eq. (12); the dashed dotted line is the prediction obtained in Dyson (1991) for a semi-infinite domain (see NoteSM for details).
The length-scale $\ell_{p}$ identified by WKB Eq. (5) plays a key role in understanding the increase of $N_{p}$ with $\alpha$ in Fig. 2a. Uncoupled oscillators spaced by $\ell$ accumulate phase differences at a rate $|g|\ell$, reaching differences $\simeq\pi$ over times $\tau_{\ell}\propto 1/|g|\ell$. Diffusive spatial couplings extend their spatial range as the square-root of time. For two points to be coupled strongly enough to lock, it is necessary that $\ell\propto\sqrt{D\tau_{\ell}}$. We conclude that the characteristic length of plateaus is $\ell_{p}$ and
$N_{p}\propto\frac{L}{\ell_{p}}=\alpha^{1/3}\,.$ (6)
The scaling (6) is consistent with each point oscillating at its own frequency ($\ell_{p}\to 0$) in the uncoupled $D\to 0$ limit and it holds for different boundary conditions, as shown in Fig. 2a. The range of validity of the scaling law (6) depends on $\beta$ in Eq. (4), which can be recast as $\beta=\left(|g|\ell_{p}/r\right)^{3}$. As $\beta$ increases for a fixed large $\alpha$ (or as $\alpha$ decreases), $N_{p}$ reduces, as confirmed in a limited range in Fig. 2a, and more globally in the $(\alpha,\beta)$ phase-space plots of Fig. 4 in Ref. NoteSM . Eventually, the minimum number of plateaus obtains (as in Fig. 2c for small $\beta$ yet moderate $\alpha$), the scaling (6) breaks down and boundary conditions matter. The strongest synchrony for Dirichlet boundary conditions has $N_{p}=2$ ; the subtler case of Neumann (no-flux) boundary conditions is considered next.
No flux entails $\nabla A\left(\pm L/2\right)=0$. Plateaus are observed again and, if $N_{p}\gg 1$, the scaling (6) still holds, demonstrating its generality (see Fig. 2a). At low $N_{p}$, the two boundary conditions differ and it appears a new state with the whole interval phase-locked, which is impossible for our Dirichlet setting. Ref. Ermentrout (1986) demonstrated locking at large $D$, and that the corresponding amplitude decreases with $|g|$ and eventually vanishes, i.e., oscillator death with a stable $A=0$ state is observed Ermentrout (1990); Reddy (2000); Ozden (2004).
Our next goal is to build on these results by studying the linear and nonlinear evolution of the dead $A=0$ and the locked states. In the linear regime, $|A|\ll 1$, the GL Eq. (1) reduces to
$\partial_{\tau}A=\partial^{2}_{\zeta}A+(r^{\prime}+i\alpha\zeta)A\,;\quad\zeta\equiv x/L\,,\,\,\,\,\tau\equiv Dt/L^{2}\,,$ (7)
where $r^{\prime}\equiv rL^{2}/D$. Eq. (7) for $r^{\prime}=0$ was introduced by Bloch Bloch (1946) and Torrey Torrey (1956) for nuclear magnetic resonance based on the following physics. A pulse ${\bm{B}}_{\perp}$ oscillating at the Larmor frequency $\omega_{L}$ is applied orthogonally to an aligning magnetic field ${\bm{B}}$. Slowly-varying ${\bm{B}}$’s are approximated by a gradient, and the linear dependence inherited by $\omega_{L}$ brings in the $\zeta$ term in Eq. (7) ; diffusion accounts for transfer of magnetization in non-homogeneous conditions. Dephasing and diffusion lead to relaxation of the transverse magnetization in the Bloch-Torrey equation, and the decay in time of its eigenmodes Dyson (1991).
The spectrum of the linear Bloch-Torrey equation is non-trivial, though, as the operator on the right-hand side of Eq. (7) is non-Hermitian. In particular, such operators often feature spectral branch points where two eigenmodes fuse Grebenkov (2022) ; Figs. 3a,b shows that is the case for the Bloch-Torrey operator Dyson (1991). For $\alpha=0$, all eigenvalues are real as the operator is Hermitian. As $\alpha$ increases, pairs of branches interconnect at branch points for successive values $\alpha_{p}$. For $\alpha>\alpha_{p}$ the corresponding pair of eigenvalues are complex conjugate, with the first branching at $\alpha_{1}\approx 18.1$ Dyson (1991) (see End Matter for details).
@@
Figure 4: Nonlinear phase-locking. The critical curve $\alpha^{4/3}\beta_{c}^{1/3}$ vs $\alpha$ that separates nonlinear locking from multiple plateaus. Convergence to $\gamma_{c}$ (dashed line) agrees with our prediction Eq. (12). Inset: The l.h.s. of Eq. (11) vs the rescaled space variable $\zeta$ (grey circles) obtained from simulations for $\alpha=2000$, $\beta=8.49\times 10^{-11}$ and our prediction (solid line) Eq. (11).
Eigenvalues $\lambda_{n}$ of the linearized GL Eq. (7) are read from the eigenvalues $\sigma_{n}$’s of the Bloch-Torrey equation ($r^{\prime}=0$) by the shift $\lambda_{n}=\sigma_{n}+r^{\prime}$ (see End Matter). While the $\sigma$’s have negative real part, the shift makes $\lambda_{n}$’s positive for $r^{\prime}>-\mathrm{Re}[\sigma_{n}]$. Fig. 3c illustrates the instabilities expected at small $\beta$’s, where the unstable drive $r$ is large (see Eq. (4)). The first pair of complex eigenvalues $\lambda_{1}^{R}\pm i\lambda_{1}^{I}$ appears at the aforementioned threshold $\alpha_{1}$. In the End Matter we show that the appearance of pairs is associated with precursors of defects, i.e., strong amplitude modulation around $x=0$ with a period $2\pi/|\lambda_{1}^{I}|$.
In the linear regime and $\lambda_{1}^{R}<0$, oscillator death takes place. For $\lambda_{1}^{R}>0$, exponential amplification takes place and nonlinear effects become relevant. The phase diagram is summarized in Fig. 3c. For small $\alpha$’s, diffusion couples the whole interval and a phase-locked state is observed Ermentrout (1986). As $\alpha$ increases, the range of diffusion shrinks, and the locked state breaks up into multiple plateaus.
While the break up in multiple plateaus was reported in Ermentrout (1986), it was concluded that “The transition from locking to plateaus remains obscure”, and the problem remained open. To solve it, we note that Fig. 3c highlights the necessity of nonlinear effects as the transition does not occur at $\alpha_{1}$, as in the linear regime. Specifically, inspection of the critical curve suggests its scaling $\beta_{c}(\alpha)\propto\alpha^{-4}$, which is the reason we introduced $\gamma$ in Eq. (4). Locking occurs for $\gamma<\gamma_{c}$ and the critical values asymptote to $\gamma_{c}\simeq 11.02$ at large $\alpha$’s (see Fig. 4).
@@
The above hints are captured and used to understand the transition as follows. Eqs. (2) and (3) are conveniently non-dimensionalized as
@@
where we have rescaled space and time as in Eq. (7), and primes denote $\zeta$-derivatives. For large $\alpha$’s, we neglect the diffusive term in Eq. (8) and assume stationarity to obtain $a\simeq\sqrt{1-\gamma\varphi^{\prime 2}}$. Inserting it into Eq. (9), we derive
$\partial_{\tau}\varphi=\zeta+\varphi^{\prime\prime}\,\frac{1-3\gamma\varphi^{\prime 2}}{1-\gamma\varphi^{\prime 2}}\,,$ (10)
which shows that nonlinearities renormalize the diffusivity. For $13\gamma\varphi^{\prime 2}>$ 1, the diffusivity turns negative and an instability is triggered. To localize it, note that $\varphi^{\prime}(-1/2)=0$ (no flux) and, if the system is stable, Eq. (10) implies $\varphi^{\prime\prime}\propto-\zeta$. It follows that the maximum of $\varphi^{\prime}$, where the negative-diffusivity instability is triggered, occurs at $\zeta=0$. Its threshold is calculated by considering the stationary Eq. (10) : $\chi^{\prime}\,\frac{1-3\gamma\chi^{2}}{1-\gamma\chi^{2}}=-\zeta$ for $\varphi^{\prime}\equiv\chi$. Integrating it between $-1/2$ and $\zeta$, we obtain

### v2

@@ Abstract
Spatial non-homogeneities can synchronize clusters of spatially-extended oscillators in different frequency plateaus. Motivated by physiological rhythms, we fully characterize the phase diagram of a Ginzburg-Landau (GL) model with a gradient of frequencies. For large gradients and diffusion, the rest state is stable, and the linear spectrum around it maps onto the non-Hermitian Bloch-Torrey equation. When complex pairs of eigenvalues turn unstable, precursors of plateaus grow, separated by defects where the GL amplitude vanishes. Nonlinear effects either saturate the amplitude of plateaus or lead to a phase-locked state, with saddle-node bifurcations separating the two regimes. In the region of plateaus, we trace the formation of defects to a non-linear renormalization of the diffusivity, and determine the scaling of their number and length vs dynamical parameters.
@@
Spatially-extended oscillations appear in a variety of situations, ranging from engineering to living, physical, and social systems Kuramoto (1984); Pikovsky (2001); Strogatz (2004). Hallmarks of nonlinearities and spatial couplings that characterize those systems are synchronization among different oscillators, and entrainment by external forcing. Synchronization is defined as the oscillation in unison of oscillators that have different frequencies yet lock because of their mutual coupling. A spectacular example is the synchronized flashing of thousands of firefly males during mating season Buck (1966); Peleg (2020). Other well-studied examples include locomotion, especially by fish Grillner (2020), and electrical waves in neural computations Ermentrout (2001); Muller (2018). Entrainment by external forcing is the locking of oscillators to the driving rhythm. A physiological instance that impacts functional magnetic resonant imaging (fMRI) is vasomotion Haddock (2005); Drew (2020). Arterioles that source blood to the brain have the diameter of their walls driven by oscillations of the surrounding smooth muscle cells. In the presence of external neuronal stimuli, a strong vasomotion increase is observed at the drive frequency Broggini (2024).
@@
A number of theoretical methods have been developed to grasp the above phenomena. Phase models Kuramoto (1984); Pikovsky (2001) are rooted in the insight that weak couplings among oscillators perturb the stable limit cycle of each oscillator in a specific manner: Transverse perturbations relax rapidly, while phase changes along the cycle can be large and do not relax (zero modes). Technically, the limit cycle and its isochrones Winfree (2001) are unperturbed, while weak perturbations drive the phase. The major simplification is that each oscillator is described by a single variable, its phase in the limit cycle, rather than the original, generally multidimensional, description (see Supplementary Material (SM) NoteSM ). The famous Kuramoto transition to synchronization Kuramoto (1984) was obtained within this framework.
@@
We focus on spatially non-homogeneous oscillatory media. Non-homogeneity may be caused by spatial variations of the parameters, e.g., the frequency of oscillators, or non-uniform external drive. Our motivation stems from two natural systems under intense experimental investigation : the aforementioned cortical vasomotion Haddock (2005); Drew (2020); Broggini (2024), and peristalsis Diamant (1969); Parsons (2020). Interstitial cells in the small intestine of mammals bear spontaneous oscillations that are synchronized by gap junction couplings, and generate rhythmic waves of contraction. Non-homogeneity is imposed by the frequencies of the oscillators in isolation decreasing aborally in a roughly linear fashion from the stomach to the colon (see Fig. 1a), as found by post hoc sectioning of the intestine Diamant (1969). In the intact intestine, couplings result in parcellation, i.e., a series of frequency plateaus as in Fig. 1a. Oscillations within the plateaus are regular, while their boundary displays “waxing and waning” (see Fig. 1b) that results from the interference of mismatched adjacent cell populations Diamant (1969).
@@
Figure 1: (Color online) Experimental peristalsis and comparison of our GL modeling. (a) Average frequencies recorded in isolated segments of cat intestine (crosses ; the black dashed line is their linear fit) and in the intact intestine (circles); (b) Time-course signals within two distinct plateaus (top and bottom), where relatively regular oscillations are observed, and in the separating region of “waxing and waning”. Data were extracted from Ref. Diamant (1969) as detailed in NoteSM . Panels (a) and (b) refer to two different experimental realizations so we do not associate the curves in (b) to specific points in (a). Panels (c) and (d) show the same curves as in (a) and (b) obtained from our GL modeling with the use of the procedure detailed in NoteSM . The plateau architecture is captured, as well as the structure of the three time courses that we locate by their colors. The inset shows the full simulation domain.
@@
To understand parcellation, it is natural to consider a chain of weakly coupled oscillators with a gradient of natural frequencies. Ref. Ermentrout (1984) pioneered this approach and demonstrated the existence of plateaus. However, as already noted in Ermentrout (1986), the strong amplitude modulation at the boundary of plateaus is incompatible with the weak coupling approximation. A first alternative is to consider chains of nonlinear oscillators, viz., van der Pol type Linkens (1974); Osipov (1998); Menzel (2011), which is also relevant for density waves in dusty plasmas Menzel (2011). Numerical simulations confirmed the existence of plateaus and evidenced “oscillator death” for large enough gradients, i.e., the presence of regions with oscillators that are unexcited BarEli (1985); Ermentrout (1990); Reddy (2000); Ozden (2004). A second alternative is to consider a Ginzburg-Landau (GL) model (see Aranson (2002)) with a gradient of natural frequencies. The advantage is that the GL equation is the universal normal form close to a supercritical Hopf bifurcation. Oscillations are described by a complex field with a phase and an amplitude, in contrast to phase models. Refs. Ermentrout (1986, 1989); Akopov (2003); Anishchenko (2005) numerically demonstrated plateaus in the GL equation and analyzed its locked states and amplitude death at large gradients (see below). In the GL formulation, the “waxing and waning” at the boundary of plateaus is associated with defects, i.e., space-time locations where the complex field vanishes and its phase is undefined Pikovsky (2001).
@@
We revisit the GL equation with a gradient of natural frequencies to characterize its phase diagram (see Fig. 2). Specifically, we consider the one-dimensional Ginzburg-Landau (GL) equation for the complex field $A(x,t)$ :
@@
which is the Stuart-Landau normal form for oscillators close to a supercritical Hopf bifucation, with diffusion accounting for spatial local coupling terms Kuramoto (1984); Pikovsky (2001). We rescaled the amplitude of $A$ to match the coefficients of the cubic and the instability term, $r$, which we take real and positive, along with the diffusivity $D$. No-flux boundary conditions entail $\nabla A\left(\pm L/2\right)=0$, the size of the domain being $L$. Note that the cubic nonlinearity is dictated by gauge invariance $A\mapsto Ae^{i\phi}$. The linear gradient $gx$ imposes non-homogeneity and breaks parity. That allows a priori for an additional term proportional to $\nabla A$ that we shall consider later, together with complex $r$ and $D$, to capture experimental data in Fig. 1.
Figure 2: Phase space portrait (a) and eigenvalues of the linear spectrum around the rest state $A=0$ (b-c). (a) Stable: The rest state is stable (and unique) in the white region at high $D$ and $g$ Ermentrout (1989). All eigenvalues have negative real part. The dashed-dotted line is the instability curve at large $\alpha$, which we predict using results for the spectrum in a semi-infinite domain Dyson (1991). Plateaus: at least a complex pair of eigenvalues has a positive real part (linear precursors of plateaus) and the nonlinear regime features multiple plateaus. Locking and nonlinear locking: A locked state Ermentrout (1986) is observed in the nonlinear regime for both regions and at least one eigenvalue of the linearized spectrum has positive real part. Their difference is that the blue region has all eigenvalues real, while the green region has at least one pair of complex eigenvalues. In other words, the blue/green regions are respectively on the left/right of the branching point shown in red in panels (b) and (c) and represented by the red dotted line in panel (a). That differentiates the approach to the locked state between the two regions (see Fig. S9). The black dashed line is the asymptotics for the fold bifurcation curve separating nonlinear locking from plateaus, as we predict by Eq. (11). (b)-(c): the real and imaginary part of the first six eigenvalues $\sigma_{n}$ vs $\alpha^{1/3}$ for the Bloch-Torrey operator defined by Eq. (4) (ordered by their real part). The imaginary part vanishes before the branching points (the first one is in red) while two complex conjugate eigenvalues per branch are present afterwards.
We start by the simplest possible solution of Eq. (1), i.e., the rest state $A=0$, which is known to be stable and unique at high values of $D$ and $g$ Ermentrout (1989). The physics is that $g$ tends to create phase differences among neighboring oscillators, while $D$ tends to homogenize their state. For strongly-driven phase differences, the only possibility left to close-up the gap in states is to shrink the limit cycle, and eventually drive it to zero BarEli (1985). To identify the term in Eq. (1) responsible for reducing the amplitude as phase differences increase, we decompose the field as $A(x,t)=a(x,t)e^{i\varphi(x,t)}$ and rewrite Eq. (1) as :
@@
which makes explicit the damping term $Da(\nabla\varphi)^{2}$.
In the linear regime around the rest state, $|A|\ll 1$, Eq. (1) reduces to
$\partial_{\tau}A=\partial^{2}_{\zeta}A+(r^{\prime}+i\alpha\zeta)A\,;\quad r^{\prime}\equiv\frac{rL^{2}}{D},\,\,\,\alpha\equiv\frac{|g|L^{3}}{D}\,$, (4)
where $\zeta\equiv x/L$ and $\tau\equiv Dt/L^{2}$. Ref. Ermentrout (1989) showed that the eigenvalues of Eq. (4) must have a negative real part if $\alpha$ is large and $r^{\prime}$ is small enough because the alternative hypothesis leads to a contradiction. An alternative path is to recognize that Eq. (4) for $r^{\prime}=0$ was introduced by Bloch Bloch (1946) and Torrey Torrey (1956) for nuclear magnetic resonance and its spectrum was characterized in Ref. Dyson (1991). The eigenvalues $\lambda_{n}$ of Eq. (4) are the Bloch-Torrey eigenvalues $\sigma_{n}$’s shifted as $\lambda_{n}=\sigma_{n}+r^{\prime}$ (see End Matter). Since the $\sigma$’s have a negative real part (see Fig. 2), the $\lambda_{n}$’s remain negative for $r^{\prime}<{\rm min}(-\mathrm{Re}[\sigma_{n}])$, in agreement with Ref. Ermentrout (1989) and the phase diagram of Fig. 2 as replotted in the variables $(\alpha,r^{\prime})$ in the supplementary Fig. S7 NoteSM .
The above alternative path brings two more results. First, it yields an analytical prediction for the instability curve at large $\alpha$’s. The difficulty with imposing the two no-flux boundary conditions in Eq. (4) is that its eigenfunctions are Airy functions Vallee (2004); Abramowitz (1965). In the limit of large $\alpha$’s, though, one can treat the domain as semi-infinite and use properties of zeros of Airy functions NoteSM ; Dyson (1991). The only dimensional combination of $g$ and $D$ (without $L$) to form an inverse time is $\left(g^{2}D\right)^{1/3}$. The upshot (see NoteSM ) for our Eq. (4) is that we predict the rest state should become unstable for $r-0.5094\left(g^{2}D\right)^{1/3}>0$, i.e., the dash-dotted line in Fig. 2 :
$\beta\equiv\frac{Dg^{2}}{r^{3}}<\beta_{c}\simeq\left(\frac{1}{0.5094}\right)^{3}\simeq 7.565\,.$ (5)
Second, the alternative path identifies branch points of the non-Hermitian operator in Eq. (4). Branch points are defined as points where two eigenmodes fuse Grebenkov (2022), as in Figs. 2b,c for the Bloch-Torrey operator Dyson (1991). For $\alpha=0$, the eigenvalues are real (the operator is Hermitian). As $\alpha$ increases, pairs of branches interconnect for successive values $\alpha_{p}$ and become complex conjugate for $\alpha>\alpha_{p}$, with the first branching at $\alpha_{1}\approx 18.1$ Dyson (1991). We show in the End Matter that conjugate pairs are associated with precursors of defects, i.e., strong amplitude modulation with a period $2\pi/|\mathrm{Im}\,\lambda_{1}|$. The first branching value $\alpha_{1}^{1/3}\simeq 2.62$ identifies the cusp at the intersection among locking, plateaus and rest states in Fig. 2.
@@
Figure 3: Nonlinear phase-locking and its transition to plateaus. (a) The critical curve $\alpha^{4/3}\beta_{c}^{1/3}$ vs $\alpha$ as identified by numerical integration of the GL equation (black points) and by the location of the saddle-node bifurcation (blue line). Inset (b): The l.h.s. of Eq. (10) vs the rescaled space variable $\zeta$ (grey circles) obtained from simulations for $\alpha=2000$, $\beta=8.50\times 10^{-11}$ and the predicted r.h.s. of Eq. (10) (solid line). (c) Bifurcation diagram of the solutions to Eqs. (7) and (8) specialized to the locked case (see text) for $\beta=1,10^{-1},10^{-2},10^{-3}$ (from darkest to lightest). Points in blue are the bifurcation points, as for the line in (a). Inset (d): zoom of the uppermost $\beta=10^{-3}$ curve.
In the unstable region (${\rm Re}\,\,\lambda_{1}>0$), exponential amplification occurs and nonlinear effects become relevant. On the left of the cusp, eigenvalues are all real and diffusion is strong enough to couple the whole interval, leading to the phase-locked state unveiled in Ref. Ermentrout (1986). The amplitude of the locked state is progressively reduced as the boundary with the stable region is approached, where it eventually vanishes Ermentrout (1986). As $\alpha$ increases, the range of diffusion shrinks, and the locked state breaks up into multiple plateaus Ermentrout (1986). To determine the nature of the change in regime, we note that Fig. 3c highlights the necessity of nonlinear effects as the transition does not occur at $\alpha_{1}$, as it does in the linear regime. Specifically, inspection of the critical curve suggests its scaling as $\beta_{c}(\alpha)\propto\alpha^{-4}$, which motivates the definition
$\gamma\equiv\left(\beta\alpha^{4}\right)^{\frac{1}{3}}=\frac{L^{4}g^{2}}{rD}\,.$ (6)
Locking occurs for $\gamma<\gamma_{c}$ and the critical values asymptote to $\gamma_{c}\simeq 11.02$ at large values of $\alpha$ (see Fig. 3a).
@@
The above hints are captured and used to understand the transition as follows. Equations (2) and (3) are conveniently non-dimensionalized as
@@
where space and time are rescaled as in Eq. (4), and primes denote $\zeta$-derivatives. For large values of $\alpha$, $a\simeq\sqrt{1-\gamma\varphi^{\prime 2}}$ and, inserting this into Eq. (8), we derive that nonlinearities renormalize the diffusivity
$\partial_{\tau}\varphi=\zeta+\varphi^{\prime\prime}\,\frac{1-3\gamma\varphi^{\prime 2}}{1-\gamma\varphi^{\prime 2}}\,.$ (9)
For $3\gamma\varphi^{\prime 2}>1$, the diffusivity turns negative and an instability is triggered at $\zeta=0$, where the maximum value of $\varphi^{\prime}$ is located. Indeed, $\varphi^{\prime}(-1/2)=0$ (no flux) and, if the system is stable, Eq. (9) implies $\varphi^{\prime\prime}\propto-\zeta$. The instability threshold is calculated in the stationary limit of Eq. (9), i.e., $\chi^{\prime}\,\frac{1-3\gamma\chi^{2}}{1-\gamma\chi^{2}}=-\zeta$ for $\varphi^{\prime}\equiv\chi$. Integrating between 1/2-1/2 and $\zeta$, we obtain